solve x if log(10^x+2)=2x+1
Answer by ikleyn(52911)

.
1.  It is seen by an unarmed eye that your writing for the equation is INCORRECT.



2.  The correct writing is this

    log(10^(x+2)) = 2x + 1.       (1)



3.  The solution is this:

    log(10^(x+2)) = x + 2.


    Hence, the equation (1) is equivalent to 

    x + 2 = 2x + 1,   or

    x = 1.


Answer.  x = 1.